# Radial Gradients and Intrinsic Scatter in MaNGA Galaxies

### v1

@@ 2 Data
In this section we explain the source and features of the data used in this work. Broadly, we use stellar population analysis with the MaNGA DR17 IFU data to extract information on age, [Fe/H], and abundance of different elements and also their corresponding gradients.
@@ 2.1 MaNGA Data
MaNGA, part of the Fourth-Generation Sloan Digital Sky Survey (SDSS IV, Blanton et al. (2017)), aims to comprehensively chart the chemical composition and kinematics of nearby galaxies. Using integral field unit (IFU) spectroscopy, MaNGA captures spectra across a large number of locations within each galaxy (Bundy et al., 2015). The observations from the IFUs (Drory et al., 2015) are fed into the BOSS spectrographs (Smee et al., 2013) mounted on the Sloan telescope (Gunn et al., 2006) at the Apogee Point Observatory. For this work, we used the latest and the final data release (DR) of the MaNGA which is DR 17 (Abdurro’uf et al., 2022). The DR 17 contains IFU observations for nearly 10000 unique galaxies at a resolution of Rsimilar-$\sim$2000 covering a wavelength range from 3000Å to 10000Å.
The MaNGA Data Analysis Pipeline (DAP) provides spatially stacked spectra in form of a datacube (Westfall et al., 2019; Law et al., 2021). It also provides stellar kinematics (V and $\sigma$), nebular emission line properties, and spectral indices (Belfiore et al., 2019). For this work, we subtracted the nebular emission lines that the DAP provided from all the corresponding spectrum before doing any further analysis.
Since we are interested in ETGs only, we first pick the galaxies which have elliptical morphology. It is well known for a fact that the elliptical galaxies have high Sérsic index value ($n$, Sérsic (1963)). Thus from the list of 10000 unique galaxies in MaNGA DR 17, we chose galaxies that have a value of $n>$2. This left us with 6273 galaxies out of 10010 unique galaxies that the MaNGA DR 17 provides. The number distribution of all these 6273 galaxies is shown in Fig. 1.
Figure 1: The distribution of all galaxies with $n>$2 as a function of velocity dispersion ($\sigma$).
@@ 2.2 Further Filtering
After choosing 6273 galaxies from MaNGA DR 17, we discarded galaxies that had Hα emission lines as we are interested only in ETGs that have very little to none star-forming activity. From the DAP datacube, if the sum of emission line flux within the Hα index band is less than 15% of the sum of observed flux within the same wavelength interval, then we select that galaxy for further analysis. This left us with 5164 galaxies. The Hα index band is defined to be 6548Å$<\lambda<$6578Å (Cohen et al., 1998). It is to be noted that the wavelength defined here is the restframe wavelength. In order to convert from observed to restframe wavelength we use:
$\lambda_{rest}=\frac{\lambda_{observed}}{(1+z)(1+\frac{v}{c})}$ (1)
In Eqn. 1, $z$ is the redshift of the galaxy, $v$ is the internal stellar velocity at the sampled location within the galaxy, and $c$ is the speed of light. We associate a global $\sigma$ to measure the stellar velocity dispersion at 0.5 Re. To estimate $\sigma$ at other annuli, we use median radial gradient results from Gonzalez (1993). We tested our analysis by cutting this mean in half and then re-deriving population parameters and found no significant change in our outcomes. However, when applying the results for $\sigma$ gradient in individual galaxies provided by Illingworth (1981), we observed substantially different results for our analysis. Given that Gonzalez (1993) offers $\sigma$ gradients for a larger sample of ETGs compared to the limited data in Illingworth (1981), we ultimately decided to adopt the results from Gonzalez (1993) for our analysis. We also could have adopted an annulus mean directly from the MaNGA DAP, but intuition drove us to use the independent mean trend to sidestep any line strength-$\sigma$ degeneracy that may exist.
@@
Fig. 2 shows the ratio of mean fluxes within the Hα index band and the continuum for the MaNGA galaxy 1-245451. The red and blue continuum are defined to be 6600Å$<\lambda<$6640Å and 6420Å$<\lambda<$6455Å respectively (Cohen et al., 1998). The average of flux within these two continuum is taken to be the mean flux within the continuum band. Since the ratio is close to one at the location of the galaxy (see Fig. 3), it shows that the algorithm that we opted to reject galaxies with Hα emission, is working optimally. All the wavelengths are converted to restframe wavelength using Eqn.1.
@@ 2.3 Creating Annuli
We divide each galaxy into 10 annuli starting from 0.1Re to 1.0Re with an increment of 0.1Re, where Re is the effective radius of the galaxy, synonymous to the half-light radius of the galaxy for Sérsic functions. The following equation is used to define the ellipse in the galaxy’s positional space.
@@
In Eqn. 2, $x$ $y$ are the RA and Dec respectively of the spaxel (spatial pixel) in consideration, $h$ $k$ are the RA and Dec respectively of the center of the galaxy, $\theta$ is the angle of inclination (East of North), $2a$ $2b$ are the semi-major and semi-minor axes respectively of the galaxy. In this equation, $a$ is used synonymously with Re. The $b$ is obtained from the $b/a$ ratio for the galaxy in question. All of $\theta$, $a$, and $b/a$ come from the Sérsic profile as provided by the MaNGA DAP. It is to be noted that this system of dividing the galaxy in several annuli is motivated from the the work of Parikh et al. (2018).
@@
Fig. 3 shows our annulus scheme. The representative spectrum for each annulus is the average spectrum for all spaxels within the same annulus. It is worth mentioning that MaNGA DAP provides masks for bad spaxels and we considered those masks while calculating the average for each annulus. The final representative spectrum is obtained the subtracting the average of the emission line spectrum from the average of the flux spectrum for each annulus. Fig. 4 shows two examples of spectra from two different regions of the galaxy. It is to be noted that we do not follow any stacking scheme. We wanted to investigate the scatter in different quantities and thus treat each galaxy individually. This possesses a problem of having very low signal-to-noise ratio in the spectrum from the peripheral region of the galaxy and also beyond 9500Å in the spectrum from any region of the galaxy. In order to mitigate this problem, we consider only up to the 7th annulus for every galaxy for any further analysis (if you consider the center part of the galaxy as 0th annulus, then up to the 6th annulus). Also in terms of element abundance analysis, we restrict ourselves to elements that do not have indices beyond 9500Å (C, N, Na, and Mg), and also any IMF analysis based on the Wing-Ford band $\lambda\lambda 9900$.
Figure 4: Spectra of MaNGA galaxy 1-245451 for innermost annulus (red) and 7th annulus (blue) from the center.
@@ 3 Analysis
In this section, we explain the pipeline for analysis used in this work. All the analyses are done on each individual spectrum from a particular annulus for a given galaxy. Unlike Parikh et al. (2018), we do not stack spectra in this work. One of the main goals for this work is to look into the intrinsic scatter in different derived quantities like the age, [Fe/H], and chemical abundances (this is discussed in details in §$\S$5.3). For this reason, we treat each galaxy individually. The problem of having lower S/N ratio is mitigated using the two strategies outlined in Section 2.3.
@@ 3.1 Spectral Analysis
Stellar population parameters are determined by fitting stellar population models to spectra. This can be achieved either by fitting the entire spectrum (Koleva et al., 2009b; Cappellari, 2017; Conroy et al., 2018). It can also be done by focusing on specific absorption line features (Trager et al., 2000a; Schiavon, 2007; Thomas et al., 2010). In this work, we adopt the latter method and concentrate on the strongest absorption features.
@@
Table 1: The spectral features used in this work. The acronyms are as follows: IB=Index Band, BC=Blue Continuum, RC=Red Continuum. The “Start” and “End” phrases are used to denote the start and end wavelengths of the passbands in Å. The “Unit” column specifies if the index used is an atomic feature (denoted by 0; units: Å) or a molecular feature (denoted by 1; units: mag).
Variable Feature IB Start IB End BC Start BC End RC Start RC End Unit
Age NaD 5876.875 5909.375 5860.625 5875.625 5922.125 5948.125 0
Hβ 4847.875 4876.625 4827.875 4847.875 4876.625 4891.625 0
Fe5270 5245.65 5285.65 5233.15 5248.15 5285.65 5318.15 0
Fe5335 5312.125 5352.125 5304.625 5315.875 5353.375 5363.375 0
[Fe/H] NaD 5876.875 5909.375 5860.625 5875.625 5922.125 5948.125 0
C24668 4634 4720.25 4611.5 4630.25 4742.75 4756.5 0
[Na/R] NaD 5876.875 5909.375 5860.625 5875.625 5922.125 5948.125 0
[C/R] C24668 4634 4720.25 4611.5 4630.25 4742.75 4756.5 0
[N/R] CN1 4142.125 4177.125 4080.125 4117.625 4244.125 4284.125 1
[Mg/R] Mg1 5069.125 5134.125 4895.125 4957.625 5301.125 5366.125 1
Mgb 5160.125 5192.625 5142.625 5161.375 5191.375 5206.375 0
@@
In §$\S$2.1 we mention the specific data products used in this work from the MaNGA DAP (Westfall et al., 2019). Currently, MaNGA DAP provides us with three different binning schemes: SPX that fits each individual spaxel with $g$-band S/N of at least 1, VOR10 is the Voronnoi binning with $g$-band S/N of at least 10 (Cappellari & Copin, 2003), and HYB10 which is a mix of both SPX and VOR10 binning schemes. We use SPX type and do not use any binning scheme. In order to find out the stellar kinematics, the data was fitted with templates from the MILESHC stellar library. The MILESHC consists of 42 stellar spectra by hierarchically-clustering the spectra in the MILES stellar library (Sánchez-Blázquez et al., 2006). We use the $\sigma$ provided by this model fit to the stellar continuum as a proxy for $\sigma$ at 0.5 Re. The emission line fitting module uses the MASTAR stellar library (Abdurro’uf et al., 2022; Yan et al., 2019). We subtracted the emission lines from the observed spectra within each IFU before going forward with further analysis.
In this work, we specifically analyzed the age, [Fe/H], and chemical abundances of C, N, Na, and Mg. Several absorption spectral features are used to determine the variables in question using the model described in Section 3.2. The strengths of the absorption features used are calculated using the Equations 2 and 3 in Worthey et al. (1994). The indices used to calculate different quantities and the corresponding passbands are given in Table 1. It is to be noted that 80 features were fed into the model while calculating our desired quantities. A full list of the 80 indices is given in Appendix B. Table 1 shows the indices that strongly affects the underlying variable. For example, CN1 feature mostly affects [N/R], but it also has mild effect on [C/R] as well. The “R” in this notation may stand for any generic heavy element, but for this work “R” can be interchangeably used with “Fe”.
@@
Figure 5: The strengths of selected features used in this work along with their respective errors for MaNGA galaxy 1-25451. Values are shown for the 1st through 7th annulus, inclusive. For graphical scaling, CN1 is multiplied by 20, and Mg1 is multiplied by 15.
@@
Fig. 5 shows the measured strengths (in Å for atomic features and in magnitude for molecular features) of major indices used in this work for the MaNGA galaxy 1-25451. For purposes of display, the original measured strengths of CN1 and its error is multiplied by 20 and same is done for Mg1 but multiplied by 15.
The errors associated with each measurement comes from the variance supplied with the MaNGA DAP data at each pixel. We started off by calculating the Standard Deviation (SD) of the weighted mean flux for both the blue and red continuum separately (continuum are defined in Table 1 for each major index). The SD in each case is square root of the inverse of the sum of inverse variances at each pixel within the continuum (blue or red). The total error from the blue and red continuum is the sum of the errors from both continuum. Next we find the fractional SD at each pixel within the index band from the variance at that pixel. Finally the error in the index strength at each pixel is calculated by using error propagation formulae applied to Equations 2 and 3 of Worthey et al. (1994). The total error is the algebraic sum of errors in each pixel within the index band. The final result is given as the SD by taking square root of this sum.
3.2 Model

### v2

@@ 2 Data
In this section, we describe the data sources and key features relevant to our analysis. Specifically, we employ stellar population modeling using the MaNGA Data Release 17 (DR17) IFU to derive stellar age, [Fe/H], elemental abundance patterns, and their corresponding spatial gradients.
@@ 2.1 MaNGA Data
The Mapping Nearby Galaxies at Apache Point Observatory (MaNGA) survey, part of the Fourth Generation Sloan Digital Sky Survey (SDSS-IV, Blanton et al. (2017)), is designed to map the internal kinematics and chemical composition of nearby galaxies. Utilizing IFU spectroscopy, MaNGA obtains spatially resolved spectra across a large number of positions within each galaxy (Bundy et al., 2015). The IFUs are connected to the BOSS spectrographs (Smee et al., 2013), which are mounted on the Sloan 2.5-meter telescope at Apache Point Observatory (Gunn et al., 2006).
For this work, we use the final and most comprehensive data release, DR17 (Abdurro’uf et al., 2022), which contains IFU spectroscopy for 10,010 unique galaxies. These spectra cover a wavelength range of 3000 Åto 10000 Åwith a spectral resolution of R $\sim$ 2000.
The MaNGA Data Analysis Pipeline (DAP) provides spatially stacked spectra in the form of datacubes (Westfall et al., 2019; Law et al., 2021), along with derived properties such as stellar kinematics (velocity and dispersion), nebular emission line fluxes, and spectral indices (Belfiore et al., 2019). For this analysis, we subtract the DAP-provided nebular emission lines from each spectrum prior to further modeling.
As our focus is on ETGs, we restrict the sample to systems with elliptical morphology. It is well established that elliptical galaxies exhibit high Sérsic index values ($n$, Sérsic (1963)). Thus, from the full set of 10,010 unique galaxies in MaNGA DR17, we select those with n>n>2, yielding a subsample of 6,273 galaxies. Fig. 2 of Domínguez Sánchez et al. (2022) and Fig. 12 of Fischer et al. (2019) show the distribution of MaNGA galaxies in terms of Sérsic index for DR17 and DR15 respectively. The number distribution of all these 6273 galaxies is shown in Fig. 1 (in orange) along with all MaNGA DR17 galaxies (in blue) and our sample with n>n>2 & without any Hα emission (in green).
Figure 1: The distribution of all galaxies from the MaNGA DR17 (blue), galaxies with n>n>2 (orange), and galaxies with n>n>2 & no Hα emission (green) as a function of velocity dispersion ($\sigma$).
@@ 2.2 Further Filtering
Following the selection of 6,273 galaxies from MaNGA DR17, we applied an additional cut to exclude systems with detectable Hα emission, as our analysis focuses exclusively on ETGs with minimal or no ongoing star formation. Specifically, we retained only those galaxies for which the integrated Hα emission line flux is less than 15% of the observed flux within the same spectral interval. Both quantities were computed within the Hα index bandpass, defined as 6548 $\lambda<$ 6578 Å(Cohen et al., 1998), using the DAP datacube products. This criterion reduced the sample to 5,164 galaxies.
All wavelengths referenced here are in the rest-frame. To convert from observed to rest-frame wavelengths, we use the following equation:
$\lambda_{\mathrm{rest}}=\frac{\lambda_{\mathrm{observed}}}{(1+z)(1+\frac{v}{c})}$ (1)
In Eqn. 1, $z$ denotes the galaxy redshift, $v$ is the internal stellar velocity at the sampled location within the galaxy, and $c$ is the speed of light. We associated a global $\sigma$ to measure the stellar velocity dispersion at 0.5 Re. To extrapolate $\sigma$ to other radial bins, we utilize median radial gradient trends derived from Gonzalez (1993). We validated this choice by reducing the gradient amplitude by 50% and re-computing the stellar population parameters; the results showed no significant deviation. However, when applying individual $\sigma$ gradient profiles from Illingworth (1981), we observed substantial differences in the derived quantities. Given that Gonzalez (1993) provides $\sigma$ gradients for a statistically larger ETG sample than the limited dataset in Illingworth (1981), we adopt the former for our analysis. Although MaNGA DAP products offer annular measurements of $\sigma$ directly, we opt instead for the external trend to mitigate potential degeneracies between line strength and $\sigma$, which may affect the robustness of abundance measurements.
@@
[…]
@@ 2.3 Creating Annuli
We divide each galaxy into 10 annular regions starting from center to 1.0Re with an increment of 0.1Re, where Re is the effective radius of the galaxy, synonymous to the half-light radius of the galaxy for Sérsic functions. The following equation is used to define the ellipse in the positional space of the galaxy.
@@
In Eqn. 2, $x$ & $y$ are the RA and Dec respectively of the spaxel (spatial pixel) in consideration, $h$ & $k$ are the RA and Dec respectively of the center of the galaxy, $\theta$ is the angle of inclination (East of North), $2a$ & $2b$ are the semi-major and semi-minor axes respectively of the galaxy. In this equation, $a$ is used synonymously with Re. The $b$ is obtained from the $b/a$ ratio for the galaxy in question. All of $\theta$, $a$, and $b/a$ come from the Sérsic profile as provided by the MaNGA DAP. It is to be noted that this system of dividing the galaxy in several annular regions is motivated from the the work of Parikh et al. (2018).
@@
Fig. 3 illustrates the annulus scheme used in this work. The representative spectrum for each annulus is the average spectrum of all spaxels within that annulus. The mean spectrum per annulus is computed by averaging the spectra of all constituent spaxels. The MaNGA DAP provides inverse variance estimates for each spaxel; we calculated the error for each annulus as the mean of these inverse variances across all spaxels within the annulus. MaNGA DAP also supplies masks for bad spaxels, which we apply when computing annular averages. The final representative spectrum for each annulus is obtained by subtracting the average emission-line spectrum from the average flux spectrum. Fig. 4 presents two example spectra from distinct galaxy regions. The top panel displays mean fluxes (solid lines) from the innermost (red) and 7th (blue) annuli alongside their corresponding emission line spectra (dashed lines). The bottom panel shows the emission line-subtracted spectra for these annuli. Shaded regions represent 1-$\sigma$ uncertainties, computed as the square root of the inverse of the DAP-provided errors (inverse variance) at each wavelength. The selected galaxies do not exhibit prominent emission line features. Notably, we do not employ any stacking scheme, as our goal is to probe scatter in measured quantities by treating each galaxy individually. This approach entails limitations. First, spectra beyond 9500Å are extremely noisy, preventing reliable index strength measurements in that region. Second, beyond the 7th or 8th annulus, flux levels decline significantly, further complicating index measurements. To mitigate these issues, we limit subsequent analyses to the 7th annulus for each galaxy (counting the galaxy center as the 0th annulus, this corresponds to up to the 6th annulus). Furthermore, in element abundance analysis, we restrict ourselves to elements with indices below 9500Å, and we omit IMF analysis based on the Wing-Ford band at $\lambda\lambda$9900.
Figure 4: Top panel: Mean flux (red/blue solid line) and emission line spectra (red/blue dashed line) from the innermost and 7th annuli for the MaNGA galaxy 1-245451. Bottom panel: Emission line subtracted spectra for the same galaxy (red: innermost annulus and blue: 7th annulus). The error shading shows the 1-$\sigma$ uncertainty in the measured fluxes.
@@ 3 Analysis
In this section, we outline the analysis pipeline used in this work. All analyses were performed on individual spectra from each annulus of a given galaxy. A key objective is to investigate the intrinsic scatter in derived properties such as age, [Fe/H], and chemical abundances (see $\S$5.3). To that end, galaxies are treated individually without stacking. The increased noise in the spectra is addressed by the strategies described in Section 2.3.
@@ 3.1 Spectral Analysis
Stellar population parameters were determined by fitting stellar population models to spectra. This can be achieved either by fitting the entire spectrum (Koleva et al., 2009b; Cappellari, 2017; Conroy et al., 2018) or by focusing on specific absorption line features (Trager et al., 2000a; Schiavon, 2007; Thomas et al., 2010). In this work, we adopt the latter method and focus on using the strongest absorption features.
@@
Table 1: The most influential index definitions used in this work. The acronyms are as follows: IB=Index Band, BC=Blue Continuum, RC=Red Continuum. “Start” and “End” denote the start or end wavelengths of the passbands in Å. The “Unit” column specifies if the accepted unit of the index is Å or magnitudes.
Parameter Feature IB Start IB End BC Start BC End RC Start RC End Unit
Age + [Fe/H] Fe4383 4369.125 4420.375 4359.125 4370.375 4442.875 4455.375 Å
Fe4531 4514.25 4559.25 4504.25 4514.25 4560.50 4579.25 Å
C24668 4634 4720.25 4611.5 4630.25 4742.75 4756.5 Å
Hβ 4847.875 4876.625 4827.875 4847.875 4876.625 4891.625 Å
Fe5270 5245.65 5285.65 5233.15 5248.15 5285.65 5318.15 Å
Fe5335 5312.125 5352.125 5304.625 5315.875 5353.375 5363.375 Å
Fe5406 5387.50 5415.00 5376.25 5387.50 5415.00 5425.00 Å
[…]
Ni3667 3655.10 3679.40 3630.40 3646.60 3685.40 3705.30 Å
[Na/R] NaD 5876.875 5909.375 5860.625 5875.625 5922.125 5948.125 Å
[Mg/R] Mg1 5069.125 5134.125 4895.125 4957.625 5301.125 5366.125 mag
Mg2 5154.125 5196.625 4895.125 4957.625 5301.125 5366.125 mag
Mg $b$ 5160.125 5192.625 5142.625 5161.375 5191.375 5206.375 Å
[N/R] CN1 4142.125 4177.125 4080.125 4117.625 4244.125 4284.125 mag
CN2 4142.125 4177.125 4083.875 4096.375 4244.125 4284.125 mag
CNO3862 3840.3 3883.4 3768.1 3812.3 3896.4 3916.2 Å
CNO4175 4129.4 4219.8 4082.5 4123.3 4243.3 4284.2 Å
[C/R] C24668 4634 4720.25 4611.5 4630.25 4742.75 4756.5 Å
Mg1 5069.125 5134.125 4895.125 4957.625 5301.125 5366.125 mag
CO4685 4626.40 4743.30 4557.30 4589.50 4805.10 4835.30 Å
CO5161 5154.3 5167.3 5108.4 5138.8 5188.4 5202.0 Å
@@
In $\S$2.1, we describe the specific data products used from the MaNGA DAP (Westfall et al., 2019). The DAP currently provides three binning schemes: SPX, which fits each individual spaxel with $g$-band S/N $\geq$ 1; VOR10, which applies Voronoi binning to achieve $g$-band S/N $\geq$ 10 (Cappellari & Copin, 2003); and HYB10, a hybrid of the two. We adopt the SPX scheme and do not apply any additional binning. Stellar kinematics are derived by fitting the data with templates from the MILESHC stellar library, which comprises 42 representative spectra constructed through hierarchical clustering of the MILES library (Sánchez-Blázquez et al., 2006). We use the $\sigma$ from this continuum fit as a proxy for velocity dispersion at 0.5 Re. Emission lines are modeled using templates from the MASTAR library (Abdurro’uf et al., 2022; Yan et al., 2019). These emission lines are subtracted from the observed IFU spectra prior to further analysis.
In this work, we specifically analyzed the light weighted age, [Fe/H] in lockstep with other heavy species, and individual chemical abundances of C, N, Na, and Mg. The reason for choosing these specific elements is mostly the ease with which the abundances of these elements can be measured. The indices corresponding to these elements are prominent and are not wrought with degeneracies. Also, we needed to keep in mind that we do not pick elements that have indices beyond 9500Å. Several absorption spectral features are used to determine the variables in question using the model described in Section 3.2. The strengths of the absorption features used are calculated using Equations 2 and 3 in Worthey et al. (1994). Selected index definitions are given in Table 1. The definitions are the standard ones from Worthey et al. (1994); Worthey & Ottaviani (1997); Serven et al. (2005). For a given annulus in a given galaxy, up to 70 out of a list of 80 indices were used to calculate age and abundance parameters. A full list of the 80 indices is given in Appendix B. Table 1 shows the indices that most strongly affect the underlying variable for a typical old galaxy. To assess this, the throw of the index is calculated and then put into units of the observational uncertainty $(\Delta I/0.3{\rm\ dex\ change\ in\ species\ }Z)/\sigma$]. The throw of any given index will of course change with observational uncertainty and with location in the age-metallicity plane, but in general the ETGs we study here could have been analyzed with only the Table 1 indices with little loss of fidelity compared to the full set. Throw values are generally between 2 and 6. Two indices appear in the list twice. Mg abundance affects Mg1 strongly, but C abundance also has an effect. C24668 is useful for measuring C abundance and also happens to be extraordinarily metallicity sensitive. The “R” in Table 1 stands for any generic heavy element. For the main results of this work, we have varied Fe in lockstep with other Fe-peak and heavier elements. In this paper, therefore, “R” can be used interchangeably with “Fe.” For example, [Fe/H] = [R/H] and [Mg/Fe] = [Mg/R].
@@
Figure 5: The strengths and uncertainties of selected features are displayed for MaNGA galaxy 1-25451. Values are shown for the 1st through 7th annulus, inclusive. For graphical scaling, CN1 is multiplied by 20, and Mg1 is multiplied by 15.
@@
Fig. 5 shows selected index strengths (in Å for atomic features and in magnitude for molecular features) of major indices used in this work for the MaNGA galaxy 1-25451. For purposes of display, the original measured strengths of CN1 and its error is multiplied by 20 and the same is done for Mg1 but multiplied by 15. A substantial number of galaxies were excluded from further analysis after several annuli produced unreliable index measurements, reducing the final sample from 5,164 to 2,968 galaxies.
The uncertainty associated with each measurement comes from the variance supplied with the MaNGA DAP data at each pixel. Under the assumption that the pixel errors are truly random, we calculate the error in the index strength using eqn. 33 of Cardiel et al. (1998).
3.2 Composite Model Grid and Parameter Estimation